2 blocks of equal mass ‘m‘ are tied together through a light string. one blocks is pulled along the line joining them with a constant force ‘F’. Find the tension in the string joining them.

Grade:9

1 Answers

Anjan
8 years ago
Net acceleration = F/2m
Also, from F.B.D. of 1 block, where block is moving with acceleration F/2m while being pulled by F and T, the net acceleration is a resultantant of the action of forces F and T. Therefore,                                                   
 F-T= m*F/2m
=> F-T= F/2 
=> T=F/2